OpenStudy (kuoministers):

Help me on this question please... I really dont know where to start thanks!!

OpenStudy (kuoministers):

\[\huge3^{x} + 5 \times 3^{x} = 54\]

OpenStudy (deadshot):

to start, we need to isolate the variables, we start by dividing both sides by 3^x, then subtract 5 from both sides, and then multiply both sides by 3^x \[3^x + 5 \times 3^x \div 3^x = 54 \div 3^x\] \[3^x + 5 - 5 = 54/3^x\]

OpenStudy (deadshot):

\[3^x \times 3^x = 54/3^x \times 3^x - 5\]

OpenStudy (deadshot):

\[9^x^2 = 54 -5\]

OpenStudy (deadshot):

\[9^x^2 = 51\] so now we find the square root of each side

OpenStudy (deadshot):

\[\sqrt{9^x^2} = \sqrt{51}\]

OpenStudy (deadshot):

the square and square root cancel each other out, so it would be \[9^x \approx 7 \]

OpenStudy (kuoministers):

im not sure but thats not the answer that is in the back of the book :P

OpenStudy (deadshot):

divide both sides by 9 \[9^x \div 9 = 7 \div 9\] \[^x = 7/9\]

OpenStudy (kuoministers):

@artix_17

OpenStudy (kuoministers):

thats not the answer in the book :P

OpenStudy (anonymous):

whats your answer?

OpenStudy (kuoministers):

2

OpenStudy (kuoministers):

i think the way he started the question was wrong cause i was told to keep the same base

OpenStudy (anonymous):

Yea, 3^x X 3^x can be written as 3^2x

OpenStudy (kuoministers):

yea i agree on that

OpenStudy (kuoministers):

\[\huge3^{2x} + 5 = 3^{3} \times 2\] this might help?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):